OpenStudy (anonymous):

Two college teams, A and B, are playing a game of tug of war. Team A applies a force of 103 newtons, but the rope does not move in any direction. What can you say about the force applied by Team B?

OpenStudy (carson889):

An object in motion will tend to stay in motion and an object at rest will tend to stay at rest , unless acted upon by an unbalanced force. Since the rope is not moving in any direction, the acceleration is thus zero. Net Force = mass * net acceleration = mass * 0 = 0. Net force = force of A - force of B. If net force = 0. Then 0 = 103 - x, where x = 103. Thus, team B must be applying an equal and opposite force of 103 newtons.

OpenStudy (anonymous):

Two college teams, A and B, are playing a game of tug of war. Team A applies a force of 103 newtons, but the rope does not move in any direction. What can you say about the force applied by Team B?